OpenStudy (anonymous):

Simplify the given expression. 7/10i

OpenStudy (anonymous):

Multiply the numerator and denominator of the expression by the conjugate of 10i which is -10i and you should get...-7i/10

OpenStudy (anonymous):

\[\frac{(7)(-10i)}{(10i)(-10i)}\]\[\frac{-70i}{-100i^2}\] Since i^2=-1 then -100*-1=100 Thus your simplified answer \[\frac{-70i}{100}\] \[\frac{-7i}{10}\]